# Collaborative Planar Pushing of Polytopic Objects with Multiple Robots in Complex Scenes

### v1

@@ Abstract
Pushing is a simple yet effective skill for robots to interact with and further change the environment. Related work has been mostly focused on utilizing it as a non-prehensile manipulation primitive for a robotic manipulator. However, it can also be beneficial for low-cost mobile robots that are not equipped with a manipulator. This work tackles the general problem of controlling a team of mobile robots to push collaboratively polytopic objects within complex obstacle-cluttered environments. It incorporates several characteristic challenges for contact-rich tasks such as the hybrid switching among different contact modes and under-actuation due to constrained contact forces. The proposed method is based on hybrid optimization over a sequence of possible modes and the associated pushing forces, where (i) a set of sufficient modes is generated with a multi-directional feasibility estimation, based on quasi-static analyses for general objects and any number of robots; (ii) a hierarchical hybrid search algorithm is designed to iteratively decompose the navigation path via arch segments and select the optimal parameterized mode; and (iii) a nonlinear model predictive controller is proposed to track the desired pushing velocities adaptively online for each robot. The proposed framework is complete under mild assumptions. Its efficiency and effectiveness are validated in high-fidelity simulations and hardware experiments. Robustness to motion and actuation uncertainties is also demonstrated.
@@ I Introduction
Non-prehensile manipulation skills such as pushing is essential when humans interact with objects, as an important complementary to prehensile skills such as stable grasping. Existing work can be found that exploits this aspect for robotic systems, mostly for a single manipulator within simple environments, e.g., Goyal et al. [8], Lynch et al. [18], Hogan and Rodriguez [11], Xue et al. [32]. However, pushing can also be beneficial for low-cost mobile robots that are not equipped with a manipulator, e.g., obstacles could be pushed away, or target objects could be pushed to destinations. As illustrated in Fig. 1, a fleet of such robots could be more efficient by collaboratively and simultaneously pushing the same object at different points with different forces. Theses combinations generate a rich set of pushing modes, e.g., long-side, short-side, diagonal, and caging, which lead to different motions of the object such translation and rotation, however with varying quality. Despite of its intuitiveness, the collaborative pushing task incorporates several challenges that are typical for contact-rich tasks, i.e., the hybrid dynamic system due to switching between different contact modes, under-actuation due to constrained contact forces, and tight kinematic and geometric constraints (such as narrow passages and sharp turns), yielding it a difficult problem not only for modeling and planning, but also for optimal control.
@@ I-A Related Work
The task and motion planning problem for prehensile manipulation is the predominant research direction with a large body of literature. It has a strong focus on the grasping policies for different objects, and the sequence of manipulating these objects for a long-term manipulation task, such as assembly in Toussaint [24], Guo and Bürger [9], re-arrangement in Kim et al. [14] and construction in Hartmann et al. [10]. On the other hand, non-prehensile planar manipulation is a challenging problem on its own for model-based planning and control without stable grasping. The single-pusher-single-slider system has been introduced as the classic model from Goyal et al. [8], where both the discrete decision for interaction modes and continuous optimization for control inputs are essential. However, the inclusion of discrete variables drastically increases the planning complexity, i.e., exponential to the planning horizon. For a small number of pre-defined modes such as sticking and sliding, this decision can be directly formulated as an integer variable in the motion planning scheme from Hogan and Rodriguez [11] thus solved via integer programs. In addition, various search methods can be applied to explore the statespace e.g., via sampling in Wang et al. [29], multi-bound tree search in Toussaint et al. [25], and backtracking in Kaelbling and Lozano-Pérez [12]. To further alleviate the complexity, human demonstrations are used in Xue et al. [32] to guide the choice of contact modes, while learning-based methods are proposed to predict the contact sequence in Simeonov et al. [21] or even raw inputs in Zeng et al. [33] for short-term tasks. However, these methods are mostly developed for a single manipulator and fix-shaped objects, of which the control and planning techniques are not applicable to multi-robot fleets.
@@
Collaborative pushing belongs to a larger application domain of multi-robot systems: cooperative object transportation, see Tuci et al. [26], which includes pushing, grasping and caging as different behaviors. The pioneer work in Kube [15] designs a multi-layer state machine for $5$ robots to push a rectangular box without direct communication in free space. A simple yet effective strategy is proposed in Chen et al. [4] to transport any convex object in obstacle-free space, i.e., the robots only push the object at positions where the direct line to the goal is occluded by the object. A combinatorial-hybrid optimization problem is formulated in Tang et al. [22] for several objects but with pre-defined pushing primitives. Fleets of heterogeneous robots are deployed in Vig and Adams [27] to clear movable obstacles via pushing, however via a pre-defined sequence of steps. Model-free approaches can be found in Wang and De Silva [28], Xiao et al. [31] that map directly the local or global scene to control inputs, which however lacks the theoretical guarantees on completeness. Other related work neglects the control aspect and focuses only on the high-level task assignment, e.g., Miyata et al. [19], García et al. [7].
@@ I-B Our Method
This work addresses the general problem of controlling a team of mobile robots to push collaboratively a polytopic object to a goal position in a complex environment, without any pre-defined contact modes or primitives. To begin with, a multi-directional feasibility analysis is conducted for a quasi-static pushing condition, given the desired object motion and a parameterized contact mode. Then, the notion of sufficient contact modes is introduced for pushing the object along a desired trajectory, based on its finite segmentation into arcs. Thus, a hierarchical hybrid search algorithm is proposed for finding the a feasible sequence of modes and the associated force profiles. It iteratively decomposes the desired trajectory via key-frames and expands the search tree via generating suitable modes between consecutive frames, in order to minimize the aforementioned feasibility loss, transition cost and control efforts. Lastly, a nonlinear model predictive controller (MPC) is proposed to track these arc segments online with the given mode, while accounting for motion uncertainties and undesired slips adaptively. Theoretical guarantees on completeness and performance are provided. Extensive high-fidelity simulations and hardware experiments are conducted to validate its efficiency and robustness.
@@ II-A Model of Workspace and Robots
Lastly, there is a single target object $\Omega\subset\widehat{\mathcal{W}}$, of which the shape is an arbitrary polygon formed by $V\geq 3$ ordered vertices ${p_{1}p_{2}\cdots p_{V}}$. Similarly, its state at time $t\geq 0$ is defined by its position and orientation, i.e. $\mathbf{x}(t)\triangleq(x,y)\in\widehat{\mathcal{W}}$ is the center of mass and $\psi\in[-\pi,\pi]$ is the rotation, $\mathbf{v}(t)\triangleq(v_{x},v_{y})$ is the linear velocity and $\omega$ is the angular velocity, and its occupied area is denoted by $\Omega(t)$. For the ease of notation, let $\mathbf{s}(t)\triangleq(x,y,\psi)\in\mathcal{S}$ denote its full state and $\mathbf{p}(t)\triangleq(v_{x},v_{y},\omega)$ its generalized velocity. In addition, its intrinsics are known a priori, including its mass $M$, the moment of inertia $\mathcal{I}$ about the center of mass, the pressure distribution at bottom surface, the coefficient of lateral friction $\mu_{c}>0$, and the coefficient of ground friction $\mu_{s}>0$.
@@ III-A2 Multi-directional Feasibility Estimation
However, due to unmodeled noises and uncertainties during the contact-rich pushing process, the target motion often deviates from the expected trajectory. Therefore, to facilitate online adaptation during execution, it is essential to estimate the feasibility in multiple directions. Denote by $\mathcal{D}$ the set of directions, which contains the desired velocity $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}$ as the main direction, and several orthogonal auxiliary directions. Since $\dim(\mathbf{p})=3$, five auxiliary directions are chosen to span the velocity space with only positive weights, i.e., $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}}\triangleq\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}}\triangleq\mathbf{e}_{3}\times\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[3]}}\triangleq\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}}\times\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[4]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[5]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[6]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[3]}}$, where $\mathbf{e}_{3}\triangleq(0,0,1)$ and $\mathcal{D}\triangleq\{1,\cdots,6\}$. Consequently, the multi-directional feasibility estimation is the weighted sum of feasibility in each direction, i.e.,
@@ III-A3 Loss Estimation for Arc Transition
Figure 4: Arc approximation (Left) and the set of allowed velocities $\mathcal{P}_{\xi}^{\scriptscriptstyle{}{\mathrm{B}}}$ (Right) visualized on a unit sphere.
@@
With Lemma 2, denote by $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}\triangleq\mathbf{Trj}(\mathbf{s}_{0},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}},\bar{t})$ the unique arc transition from $\mathbf{s}_{0}$ to $\mathbf{s}_{\texttt{G}}$, of which the arc length is given by $\|\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}\|\triangleq\|\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}\bar{t}\|_{2}$. Consequently, a loss estimation is proposed for choosing a mode $\xi$ under an arc transition $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}$:
@@ III-A4 Mode Generation for Arc Transition
With the proposed loss function in (12), the optimal mode $\xi^{\star}$ for a desired arc transition $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}$ can be determined, e.g., by selecting from a finite set of possible modes as in Tang et al. [22], Wang and De Silva [28]. However, as described earlier, the set of all possible modes $\Xi$ in this work is infinite, of which the parameter space grows exponentially with the number of robots and the boundary length of the target. To tackle this issue, a fast and scalable method to generate suitable modes for a given arc transition is proposed, called mode generation via sparse optimization (MG-SO). To begin with, each side $(p_{v},p_{v+1})$ of the target is divided into $N_{v}>0$ segments, $\forall v\in\{1,\cdots,V\}$. The center of each segment is denoted by $p_{v,i}$, $\forall i\in\{1,\cdots,N_{v}\}$. Then, consider a virtual mode $\xi^{\star}$ that employs $\{p_{v,i},\forall i,\forall v\}$ as the contact points, i.e., $\xi^{\star}\triangleq p_{\scriptscriptstyle{}{1,1}}\cdots p_{\scriptscriptstyle{}{1,N_{1}}}\cdots p_{\scriptscriptstyle{}{V,1}}\cdots p_{\scriptscriptstyle{}{V,N_{V}}}$. Note that the total number of contact points is $N_{\mathcal{V}}=\sum_{v=1}^{V}N_{v}$, which is much larger than the number of robots $N$. The associated force $\mathbf{F}_{\xi^{\star}}\triangleq(\mathbf{F}_{\texttt{n}},\,\mathbf{F}_{\texttt{t}})$ follows the definition in (2), among which the forces along the multiple directions $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[d]}}$ in (10) are given by $\mathbf{F}_{\xi^{\star}}^{\scriptscriptstyle{}{[d]}}\triangleq(\mathbf{F}^{\scriptscriptstyle{}{[d]}}_{\texttt{n}},\,\mathbf{F}^{\scriptscriptstyle{}{[d]}}_{\texttt{t}})$, $\forall d\in\mathcal{D}$. Subsequently, the following sparse optimization problem is formulated:

### v2

@@ Abstract
Pushing is a simple yet effective skill for robots to interact with and further change the environment. Related work has been mostly focused on utilizing it as a non-prehensile manipulation primitive for a robotic manipulator. However, it can also be beneficial for low-cost mobile robots that are not equipped with a manipulator. This work tackles the general problem of controlling a team of mobile robots to push collaboratively polytopic objects within complex obstacle-cluttered environments. It incorporates several characteristic challenges for contact-rich tasks such as the hybrid switching among different contact modes and under-actuation due to constrained contact forces. The proposed method is based on hybrid optimization over a sequence of possible modes and the associated pushing forces, where (i) a set of sufficient modes is generated with a multi-directional feasibility estimation, based on quasi-static analyses for general objects and any number of robots; (ii) a hierarchical hybrid search algorithm is designed to iteratively decompose the navigation path via arc segments and select the optimal parameterized mode; and (iii) a nonlinear model predictive controller is proposed to track the desired pushing velocities adaptively online for each robot. The proposed framework is complete under mild assumptions. Its efficiency and effectiveness are validated in high-fidelity simulations and hardware experiments. Robustness to motion and actuation uncertainties is also demonstrated. More examples and videos are available at https://zilitang.github.io/Collaborative-Pushing.
@@ I Introduction
Non-prehensile manipulation skills such as pushing are essential when humans interact with objects, as an important complementary to prehensile skills such as stable grasping. Existing work can be found that exploits this aspect for robotic systems, mostly for a single manipulator within simple environments, e.g., Goyal et al. [8], Lynch et al. [19], Hogan and Rodriguez [11], Xue et al. [32]. However, pushing can also be beneficial for low-cost mobile robots that are not equipped with a manipulator, e.g., obstacles could be pushed away, or target objects could be pushed to destinations. As illustrated in Fig. 1, a fleet of such robots could be more efficient by collaboratively and simultaneously pushing the same object at different points with different forces. Theses combinations generate a rich set of pushing modes, e.g., long-side, short-side, diagonal, and caging, which lead to different motions of the object such translation and rotation, however with varying quality. Despite of its intuitiveness, the collaborative pushing task incorporates several challenges that are typical for contact-rich tasks, i.e., the hybrid dynamic system due to switching between different contact modes, under-actuation due to constrained contact forces, and tight kinematic and geometric constraints (such as narrow passages and sharp turns), yielding it a difficult problem not only for modeling and planning, but also for optimal control.
@@ I-A Related Work
The task and motion planning problem for prehensile manipulation is the predominant research direction with a large body of literature. It has a strong focus on the grasping policies for different objects, and the sequence of manipulating these objects for a long-term manipulation task, such as assembly in Toussaint [24], Guo and Bürger [9], re-arrangement in Kim et al. [14] and construction in Hartmann et al. [10]. On the other hand, non-prehensile planar manipulation is a challenging problem on its own for model-based planning and control without stable grasping. The single-pusher-single-slider system has been introduced as the classic model from Goyal et al. [8], where both the discrete decision for interaction modes and continuous optimization for control inputs are essential. However, the inclusion of discrete variables drastically increases the planning complexity, i.e., exponential to the planning horizon. For a small number of pre-defined modes such as sticking and sliding, this decision can be directly formulated as an integer variable in the motion planning scheme from Hogan and Rodriguez [11] thus solved via integer programs. In addition, various search methods can be applied to explore the statespace e.g., via sampling in Wang et al. [29], multi-bound tree search in Toussaint et al. [25], and backtracking in Kaelbling and Lozano-Pérez [12]. To further alleviate the complexity, human demonstrations are used in Xue et al. [32], Le et al. [17] to guide the choice of contact modes, while learning-based methods are proposed to predict the contact sequence in Simeonov et al. [21] or even raw inputs in Zeng et al. [33] for short-term tasks. However, these methods are mostly developed for a single manipulator and fix-shaped objects, of which the control and planning techniques are not applicable to multi-robot fleets.
@@
Collaborative pushing belongs to a larger application domain of multi-robot systems: cooperative object transportation, see Tuci et al. [26], which includes pushing, grasping and caging as different behaviors. The pioneer work in Kube [15] designs a multi-layer state machine for $5$ robots to push a rectangular box without direct communication in free space. A simple yet effective strategy is proposed in Chen et al. [4] to transport any convex object in obstacle-free space, i.e., the robots only push the object at positions where the direct line to the goal is occluded by the object. A combinatorial-hybrid optimization problem is formulated in Tang et al. [22] for several objects but with pre-defined pushing primitives. Fleets of heterogeneous robots are deployed in Vig and Adams [27] to clear movable obstacles via pushing, however via a pre-defined sequence of steps. Model-free approaches can be found in Xiao et al. [31] that map directly the scene to control inputs but lacks the theoretical guarantees on completeness. Other related work neglects the control aspect and focuses only on the high-level task assignment, e.g., García et al. [7].
@@ I-B Our Method
This work addresses the general problem of controlling a team of mobile robots to push collaboratively a polytopic object to a goal position in a complex environment, without any pre-defined contact modes or primitives. To begin with, a multi-directional feasibility analysis is conducted for a quasi-static pushing condition, given the desired object motion and a parameterized contact mode. Then, the notion of sufficient contact modes is introduced for pushing the object along a desired trajectory, based on its finite segmentation into arcs. Thus, a hierarchical hybrid search algorithm is proposed for finding a feasible sequence of modes and the associated force profiles. It iteratively decomposes the desired trajectory via key-frames and expands the search tree via generating suitable modes between consecutive frames, in order to minimize the aforementioned feasibility loss, transition cost and control efforts. Lastly, a nonlinear model predictive controller (MPC) is proposed to track these arc segments online with the given mode, while accounting for motion uncertainties and undesired slips adaptively. Theoretical guarantees on completeness and performance are provided. Extensive high-fidelity simulations and hardware experiments are conducted to validate its efficiency and robustness.
@@ II-A Model of Workspace and Robots
Lastly, there is a single target object $\Omega\subset\widehat{\mathcal{W}}$, of which the shape is an arbitrary polygon formed by $V\geq 3$ ordered vertices ${p_{1}p_{2}\cdots p_{V}}$. Similarly, its state at time $t\geq 0$ is defined by its position and orientation, i.e. $\mathbf{x}(t)\triangleq(x,y)\in\widehat{\mathcal{W}}$ is the center of mass and $\psi\in[-\pi,\pi]$ is the rotation, $\mathbf{v}(t)\triangleq(v_{x},v_{y})$ is the linear velocity and $\omega$ is the angular velocity, and its occupied area is denoted by $\Omega(t)$. For the ease of notation, let $\mathbf{s}(t)\triangleq(x,y,\psi)\in\mathcal{S}$ denote its full state and $\mathbf{p}(t)\triangleq(v_{x},v_{y},\omega)$ denote its generalized velocity. In addition, its intrinsics are known a priori, including its mass $M$, the moment of inertia $\mathcal{I}$ about the center of mass, the pressure distribution at bottom surface, the coefficient of lateral friction $\mu_{c}>0$, and the coefficient of ground friction $\mu_{s}>0$.
@@ III-A2 Multi-directional Feasibility Estimation
However, due to unmodeled noises and uncertainties during the contact-rich pushing process, the target motion may deviates from the expected trajectory. Therefore, to facilitate online adaptation during execution, it is essential to estimate the feasibility in multiple directions. Denote by $\mathcal{D}$ the set of directions, which contains the desired velocity $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}$ as the main direction, and several orthogonal auxiliary directions. Since $\dim(\mathbf{p})=3$, five auxiliary directions are chosen to span the velocity space with only positive weights, i.e., $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}}\triangleq\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}}\triangleq\mathbf{e}_{3}\times\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[3]}}\triangleq\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}}\times\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[4]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[1]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[5]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[2]}},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[6]}}\triangleq-\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[3]}}$, where $\mathbf{e}_{3}\triangleq(0,0,1)$ and $\mathcal{D}\triangleq\{1,\cdots,6\}$. Consequently, the multi-directional feasibility estimation is the weighted sum of feasibility in each direction, i.e.,
@@ III-A3 Loss Estimation for Arc Transition
Figure 4: Arc transitions (Left) and the set of allowed velocities $\mathcal{P}_{\xi}^{\scriptscriptstyle{}{\mathrm{B}}}$ (Right) visualized on a unit sphere.
@@
With Lemma 2, we denote by $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}\triangleq\mathbf{Trj}(\mathbf{s}_{0},\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}},\bar{t})$ the unique arc transition from $\mathbf{s}_{0}$ to $\mathbf{s}_{\texttt{G}}$, of which the arc length is given by $\|\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}\|\triangleq\|\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}\bar{t}\|_{2}$. Consequently, a loss estimation is proposed for choosing a mode $\xi$ under an arc transition $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}$:
@@ III-A4 Mode Generation for Arc Transition
With the proposed loss function in (12), the optimal mode $\xi^{\star}$ for a desired arc transition $\overgroup{\mathbf{s}_{0}\mathbf{s}_{\texttt{G}}}$ can be determined. However, as described earlier, the set of all possible modes $\Xi$ in this work is infinite, of which the parameter space grows exponentially with the number of robots and the boundary length of the target. To tackle this issue, a fast and scalable method to generate suitable modes for a given arc transition is proposed, called mode generation via sparse optimization (MG-SO). To begin with, each side $(p_{v},p_{v+1})$ of the target is divided into $N_{v}>0$ segments, $\forall v\in\{1,\cdots,V\}$. The center of each segment is denoted by $p_{v,i}$, $\forall i\in\{1,\cdots,N_{v}\}$. Then, consider a virtual mode $\xi^{\star}$ that employs $\{p_{v,i},\forall i,\forall v\}$ as the contact points, i.e., $\xi^{\star}\triangleq p_{\scriptscriptstyle{}{1,1}}\cdots p_{\scriptscriptstyle{}{1,N_{1}}}\cdots p_{\scriptscriptstyle{}{V,1}}\cdots p_{\scriptscriptstyle{}{V,N_{V}}}$. Note that the total number of contact points is $N_{\mathcal{V}}=\sum_{v=1}^{V}N_{v}$, which is much larger than the number of robots $N$. The associated force $\mathbf{F}_{\xi^{\star}}\triangleq(\mathbf{F}_{\texttt{n}},\,\mathbf{F}_{\texttt{t}})$ follows the definition in (2), among which the forces along the multiple directions $\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}_{\scriptscriptstyle{}{[d]}}$ in (10) are given by $\mathbf{F}_{\xi^{\star}}^{\scriptscriptstyle{}{[d]}}\triangleq(\mathbf{F}^{\scriptscriptstyle{}{[d]}}_{\texttt{n}},\,\mathbf{F}^{\scriptscriptstyle{}{[d]}}_{\texttt{t}})$, $\forall d\in\mathcal{D}$. Subsequently, the following sparse optimization problem is formulated: